40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 283 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  2638   Wed Feb 24 16:11:15 2010 JenneUpdatePEMSeismic witnesses near MC1 tank moved

Since we're going to open the MC1 tank tomorrow, I've moved the MC1 accelerometers and the Guralp over to underneath MC2 for the vent.  I'll reconnect them later.

  2637   Wed Feb 24 12:08:31 2010 KojiUpdateComputersRFM goes red -> recovered by the nuclear option

Most of the RFM went red this morning. I took the nuclear option and it seemed to be recovered.

  2636   Tue Feb 23 19:07:43 2010 kiwamu and steveConfigurationVACvent is completed

Quote:

Kiwamu and Steve have started venting the 40m vacuum envelope.

Preparation:

centered oplevs at resonating cavities,

ITM references were set by green pointer from the ends by Koji,

closed  PSL shutter and placed manual block into beam path,

checked  jamnuts in locked positions on bellows,

turned  HV off at PZT-Jena "steering mirror" power supply and OMC HV ps

checked  particle counts,

switched oplev servos off,

set up N2 cylinder to start vent from 1e-6 Torr to 25 Torr,

have ~ 6 cylinders of instrument grade compressed air to bring envelope from 25 Torr to 760 Torr

 

All three cranes were wiped off today.

 

 Kiwamu has completed the vent.

Attachment 1: vent2010feb23.jpg
vent2010feb23.jpg
  2635   Tue Feb 23 19:00:45 2010 kiwamuConfigurationVACvent finished

The vent has been finished.

Now the pressure inside the chamber is 760 torr, and it's getting equilibrium with the atmospheric pressure.

Therefore we are ready and can open the door of the chamber tomorrow.

  2634   Tue Feb 23 16:42:02 2010 ranaConfigurationComputer Scripts / ProgramsSVN restarted on NODUS

I ran the start Apache script as described by Yoichi in the WIki. SVN back up.

  2633   Tue Feb 23 16:20:44 2010 steveConfigurationVACoutgassing + leak

Quote:

As preparation for the upcoming planned power outage we turned turbos, RGA off and closed valves.

IFO chamber is not pumped now. Small leaks and out gassing will push the pressure up slowly. At 3 mTorr of P1 the PSL output shutter

will be closed by the interlock.

It is OK to use light in the IFO up to this point.

 Our vacuum was not pumped for 4 days. The computers were not up when we started pumping again. The manual reading on P1 was 15 mTorr.

This means that our outgassing plus leakrate is  ~3.8 mTorr /day 

3-5 mTorr / day is normal

Attachment 1: outgleak.jpg
outgleak.jpg
  2632   Tue Feb 23 14:56:24 2010 steveUpdateGeneralETMX optical table mass

Our janitor dropped one 48" long  fluorescent tube on the top cover of ETMX-isct. This accident made glasses fly all over the place.

He cleaned up nicely, but please beware of small glass pieces around ETMX chamber.

We did not clean up on the table in order to reserve oplev and green ITMX pointing alignment.

Plot below shows that the alignment was not effected.

 

 

Attachment 1: etmxmass.jpg
etmxmass.jpg
  2631   Tue Feb 23 13:37:04 2010 kiwamu and steveConfigurationVACventing the 40m vac envelope

Kiwamu and Steve have started venting the 40m vacuum envelope.

Preparation:

centered oplevs at resonating cavities,

ITM references were set by green pointer from the ends by Koji,

closed  PSL shutter and placed manual block into beam path,

checked  jamnuts in locked positions on bellows,

turned  HV off at PZT-Jena "steering mirror" power supply and OMC HV ps

checked  particle counts,

switched oplev servos off,

set up N2 cylinder to start vent from 1e-6 Torr to 25 Torr,

have ~ 6 cylinders of instrument grade compressed air to bring envelope from 25 Torr to 760 Torr

 

All three cranes were wiped off today.

 

  2630   Tue Feb 23 06:47:57 2010 KojiUpdateGeneralIFO situations / low power MC lock

Work on 22nd Monday:

[MC recovery]

- Tried to lock MC after the computer recovery by Joe.
- A lot of higher modes. I can touch the input periscope or the MC mirrors.
- First tried to align the MC mirrors. MC1 was aligned against the MC REFL PD. MC2/3 was aligned to maximize the transmitted power.
- After the alignment, I got the MC Trans Sum ~8V. Also I saw the flashing of the arm cavities. I decided to take this alignment although the beam looks little bit clipped by the faraday.

[IFO alignment recovery]

- Aligned the arms for TEM00 manually.
- Arm alignment script seems not working now. This could come from the move of the end QPDs
- PRMI/DRMI were aligned. All alignment values saved.

[Low power MC]

[Optical config]

- I fixed the MCT CCD camera. It is quite useful to align the MC.

- Inserted HWP+Cube PBS+HWP combo in the MC incident path.
- First HWP and PBS adjust the light power. The second HWP is fixed at 342deg such that it restores the poralization to S.
- The incident power was measured by the SCIENTECH power meter. Offset of 3mW was subtracted in the table below.

HWP1 angle P_MC_incident comment
126deg 1.03W Max
100 0.39  
90 0.098 Low power max
85 0.021 Low power nominal

- HWP1 85deg is the nominal.

- I needed to touch the steering mirror (indicated by the picture) to obtain TEM00.
  The alignment of the HWPs and the cube PBS didn't change the mode. Thermal lense of the cube?

- I could not lock the MC with the incident power below 100mW. So the BS in the MC REFL path was replaced by a total reflector (Y1-45S).
- This increased the power on the MC REFL PD x10 of the previous. NOW WE ARE CONSTRAINED BETWEEN 81deg~90deg. DON'T ROTATE FURTHER!
- The original BS was stored on the AP table as shown in the picture.
- This total reflector disabled the MC WFS QPDs. We can't use them.

[Lock of the MC with 20mW incident]
- Disable the MC autolocker.
- Disable the MC WFS.
- Run
  /cvs/cds/caltech/scripts/MC/mcloopson
- Turn on the MCL servo.
- Set the MCL gain to 1.5 (it was nominally 0.3 for the high power)
- Just wait until lock.

[Gain boost after the lock] ...If you like to have more gain
- There was almost no room to increase the MCL gain.
- MC_REFL_GAIN can be increased from +6dB to +20dB
  ezcawrite "C1:IOO-MC_REFL_GAIN" 20
- MC_VCO_GAIN can be increased from -3dB to +2dB
  ezcawrite "C1:IOO-MC_VCO_GAIN" 2
- Crank the FSS gains
  ezcawrite C1:PSL-FSS_MGAIN `ezcaread -n C1:PSL-STAT_FSS_NOM_C_GAIN`
  ezcawrite C1:PSL-FSS_FASTGAIN `ezcaread -n C1:PSL-STAT_FSS_NOM_F_GAIN`

[If lock is lost]
- Run
  /cvs/cds/caltech/scripts/MC/mcdown

Attachment 1: MC_incident.png
MC_incident.png
Attachment 2: MC_REFL.png
MC_REFL.png
  2629   Mon Feb 22 21:07:26 2010 JenneUpdateCOCMagnets glued to ITMX

[Kiwamu, Jenne]

The magnets + dumbbell standoffs have been glued to ITMX.  We're waiting overnight for them to dry. 

Since I broke one of the magnet + dumbbells on the ITMY set, we've glued another dumbbell to the 6th magnet, and it should be ready for us to glue to ITMY tomorrow, once ITMX is dry and out of the fixture.  This doesn't put us behind schedule at all, so that's good.

We had been concerned that there might be a problem with the arm of the guiderod fixture being glued to ITMY, but it was fine after all.  Everything is going smoothly so far.

 

[Zach, Mott]

Zach and Mott are almost prepared to start cutting the viton for the earthquake stops.  We need 2 full sets by Wednesday morning, when we expect to begin hanging the ITMs.

  2628   Mon Feb 22 13:08:27 2010 josephbUpdateComputersMinor tweaks to c1omc

While working on c1omc, I created a .cshrc file in the controls home directory, and had it source the cshrc.40m file so that useful shortcuts like "target" and "c" work, among other things.  I also fixed the resolv.conf file so that it correctly uses linux1 as its name server (speeding up ssh login times).

  2627   Mon Feb 22 12:48:31 2010 josephb, alex, kojiUpdateComputersFE machines now coming up

Even after bringing up the fb40m, I was unable to get the front ends to come up, as they would error out with an RFM problem.

We proceeded to reboot everything I could get my hands on, although its likely it was daqawg and daqctrl which were the issue, as on the C0DAQ_DETAIL screen their status had been showing as 0xbad, but after the reboot showed up as 0x0.  They had originally come up before the frame builder had been fixed, so this might have been the culprit.  In the course of rebooting, I also found c1omc and c1lsc had been turned off as well, and turned them on.

After this set of reboots, we're now able to bring the front ends up one by one.

  2626   Mon Feb 22 11:46:55 2010 josephbUpdateComputersfb40m

I fixed the JetStor 416S raid array IP address by plugging in my laptop to its ethernet port, setting my IP to be on the same subnet, and using the web interface.  (After finally tracking down the password, it has been placed in the usual place).

After this change, I powered up the fb40m2 machine and reboot the fb40m machine. This seems to have made all the associated lights green.

Data viewer is working such that is recording from the point I fixed the JetStor raid array and did the fb40m reboot.  It also can go back in time before the IP switch over.

  2625   Mon Feb 22 11:42:48 2010 KojiUpdateGeneralPrep for Power Supply Stop

Turned on the power supply for the oplev lasers.
Turned on the power of the aux NPRO.
Turned on some of the Sorensen at 1X1.
Fixed the thermal output to round -4.0.
Locked PMC / MZ.

Waiting for the computers recovering.

  2624   Mon Feb 22 11:38:05 2010 joe, jenne and steveConfigurationVACvacuum is back to normal

Morning condition: vacuum rack power is still off, no MEDM screen reading.....meaning unknown vacuum pressure.We closed PSL shutter immediately.

Joe restored c1iscepis and Jenne powered up the vac-rack UPS. Now the rest of the vac-rack power were restored from starting at the top to bottom.

P1 was reading 15 mTorr.  We restarted pumps and  set vacuum valve positions. V1 opening required Rob's recipe of elog # 1863 to defeat interlock that

has a non communicating gauge: PTP1

CC1 pressure just reached 1e-6 Torr at VAC NORMAL configuration.

  2623   Mon Feb 22 10:25:37 2010 JenneUpdateCOCITMY standoff and guiderod epoxied

This work happened on Friday, after Nodus and the elog went down....

[Jenne, Kiwamu]

The guiderod and standoff for ITMY were epoxied, and left drying over the weekend on the flow bench under a foil tent.  The flow bench was off for the weekend, so we made tents which hopefully didn't have any place for dust to get in and settle on the mirrors.

There is a small chance that there will be a problem with glue on the arm of the fixture holding the guiderod to the optic.  Kiwamu and I examined it, and hopefully it won't stick.  We'll check it out this afternoon when we start getting ready for gluing magnets onto optics this afternoon.

  2622   Mon Feb 22 09:45:34 2010 josephbUpdateGeneralPrep for Power Supply Stop

Quote:

Autoburts have not been working since the network changeover last Thursday.

Last snapshot was around noon on Feb 11...  


It turns out this happened when the IP address got switched from 131.... to 192.... Here's the horrible little piece of perl code which was failing:

$command = "/usr/sbin/ifconfig -a > $temp";
   system($command);

   open(TEMP,$temp) || die "Cannot open file $temp\n";
   $site = "undefined";
   #                                                                                                     
   # this is a horrible way to determine site location                                                   
   while ($line = <TEMP>) {
     if ($line =~ /10\.1\./) {
       $site = "lho";
     } elsif ($line =~ /10\.100\./) {
       $site = "llo";
     } elsif ($line =~ /192\.168\./) {
       $site = "40m";
     }
   }
   if ($site eq "undefined") {
     die "Cannot Determine Which LIGO Observatory this is\n";

I've now put in the correct numbers for the 40m...and its now working as before. I also re-remembered how the autoburt works:

1) op340m has a line in its crontab to run /cvs/cds/caltech/burt/autoburt/burt.cron (I've changed this to now run at 7 minutes after the hour instead of at the start of the hour).

2) burt.cron runs /cvs/cds/scripts/autoburt.pl (it was using a perl from 1999 to run this - I've now changed it to use the perl 5.8 from 2002 which was already in the path).

3) autoburt.pl looks through every directory in 'target' and tries to do a burt of its .req file.

Oh, and it looks like Joe has fixed the bug where only op440m could ssh into op340m by editing the host.allow or host.deny file (+1 point for Joe).

But he forgot to elog it (-1 point for Joe).®

I knew there was going to be a script somewhere with a hard coded IP address.  My fault for missing it.  However, in regards to the removal of op340m's host.deny file, I did elog it here.  Item number 5.

  2621   Mon Feb 22 07:25:58 2010 ranaUpdateGeneralPrep for Power Supply Stop

Autoburts have not been working since the network changeover last Thursday.

Last snapshot was around noon on Feb 11...  


It turns out this happened when the IP address got switched from 131.... to 192.... Here's the horrible little piece of perl code which was failing:

$command = "/usr/sbin/ifconfig -a > $temp";
   system($command);

   open(TEMP,$temp) || die "Cannot open file $temp\n";
   $site = "undefined";
   #                                                                                                     
   # this is a horrible way to determine site location                                                   
   while ($line = <TEMP>) {
     if ($line =~ /10\.1\./) {
       $site = "lho";
     } elsif ($line =~ /10\.100\./) {
       $site = "llo";
     } elsif ($line =~ /192\.168\./) {
       $site = "40m";
     }
   }
   if ($site eq "undefined") {
     die "Cannot Determine Which LIGO Observatory this is\n";

I've now put in the correct numbers for the 40m...and its now working as before. I also re-remembered how the autoburt works:

1) op340m has a line in its crontab to run /cvs/cds/caltech/burt/autoburt/burt.cron (I've changed this to now run at 7 minutes after the hour instead of at the start of the hour).

2) burt.cron runs /cvs/cds/scripts/autoburt.pl (it was using a perl from 1999 to run this - I've now changed it to use the perl 5.8 from 2002 which was already in the path).

3) autoburt.pl looks through every directory in 'target' and tries to do a burt of its .req file.

Oh, and it looks like Joe has fixed the bug where only op440m could ssh into op340m by editing the host.allow or host.deny file (+1 point for Joe).

But he forgot to elog it (-1 point for Joe).®

  2620   Sun Feb 21 17:44:35 2010 ranaUpdateGeneralPrep for Power Supply Stop
  • Turned on the RAID attached to linux1 (its our /cvs/cds disk)
  • Turned on linux1 (it needed a keyboard and monitor in order to be happy - no fsck required)
  • Turned on nodus (and started ELOG) + all the control room machines
  • Turned on B/W monitors
  • Untaped fridge


  • Found several things OFF which were not listed in the Wiki...
  • Turned ON the 2 big isolation transformers (next to Steve's desk and under the printer). These supply all of the CDS racks inside.
  • ~75% of the power strips were OFF in the CDS racks ?? I turned on as many as I could find (except the OMC).
  • Switched on and keyed on all of the FE and SLOW crates in no particular order. Some of the fans sound bad, but otherwise OK.
  • Turned on all of the Sorensens that are labeled.
  • Turned ON the linear supplies close to the LSC rack.
  • ON the Marconis - set them according to the labels on them (probably out-dated).
  • After restoring power to the PSL enclosure (via the Isolation Transformer under the printer) turned the Variac ON and HEPA on full speed.
  • Plugged in the PSs for the video quads. Restored the Video MUX settings - looks like we forgot to save the correct settings for this guy...


PSL


1) Turned on the chiller, then the MOPA, then the RC's Heater power supply.
2) Shutter is open, laser is lasing, PMC is locked.
3) RC temperature is slowly rising. Will probably be thermalized by tomorrow.

Sun Feb 21 20:04:17 2010
Framebuilder is not mounting its RAID frames - in fact, it doesn't mount anything because the mountall command is failing on the RAID with the frames. The Jetstor RAID is also not responding to ping. Looks like the JetStor RAID which has all of our frames is still on the old 131 network, Joe.
  2619   Fri Feb 19 16:40:43 2010 kiwamuUpdateGreen Lockingrearrange the optics on the end table

Koji and kiwamu

The existing optics on the ETMX/ETMY end table were rearranged in this morning.

 


The main things we have done are -

1. relocation of the optical levers for ETMs ( as mentioned in koji's entry )

This relocation can make a space so that we can setup the green locking stuffs.

The optical path of the green locking is planed to start from the right top corner on the table, therefore we had to relocate the oplevs toward the center of the table.

2. relocation of the lens just before the tube

Because we are going to shoot the green beam into the arm cavity, we don't want to have any undesired lenses before the cavity.

For this reason we changed the position of the lens, it was standing just in front of the tube, now it's standing on the left side of the big mirror standing center top.

Since we did not find a significant change in its the spot size of the transmitted light, we did not change the position of all the TRANS_MON_PDs and its mirrors. And they are now well aligned.

Attachment1: ETMX end table

Attachment2: ETMY end table

Attachment 1: DSC_1202.JPG
DSC_1202.JPG
Attachment 2: DSC_1207.JPG
DSC_1207.JPG
  2618   Fri Feb 19 15:29:14 2010 kiwamuUpdateCOCGluing dumbbells and magnets

Jenne and kiwamu

We have glued the dumbbells to the magnets that will be used for the ITMs

We made two sets of glued pair of the dumbbell and the magnet ( one set means 6 pairs of the dumbbell and the magnet. Therefore in total we got 12 pairs. )

You can see the detailed procedure we did on the LIGO document E990196.

Actually we performed one different thing from the documented procedure;

we made scratch lines on the surface of the both dumbbells and magnets by a razor blade.

According to Steve and Bod, these scratch make the strength of the glues stronger.

Now the dumbbell-magnet pairs are on the flow bench in the clean room, and supported by a fixture Betsy sent us.

 

- -  notes

On the bench the left set is composed by magnets of 244 +/- 3 Gauss and the right set is 255 +/- 3 Gauss.

 

  2617   Fri Feb 19 13:28:44 2010 KojiUpdateGeneralPrep for Power Supply Stop

- ETMX/ETMY oplev paths renewed. The nominal gain for ETMY YAW was reversed as a steering mirror has been put.
- Oplevs/QPDs cenrtered except for the MCT QPD.
- SUS snapshots updated
- QPD/Aligment screenshots taken

40m Wiki: Preparation for power supply stop

Attachment 1: screen_shot.png
screen_shot.png
  2616   Fri Feb 19 10:18:19 2010 JenneUpdateVACThe P1 vac pressure is almost to 3mTorr

The Vac pressure measured at P1 is at 2.5mTorr.  I expect we'll hit 3mTorr sometime this afternoon, at which point (according to Steve) the interlock will shut the shutter, and we won't have light in the IFO.  Anything which needs to happen with light in the IFO before Monday needs to happen fairly soon.

Attachment 1: VacPressureAlmostShutoffLaser_19Feb2010.png
VacPressureAlmostShutoffLaser_19Feb2010.png
  2615   Fri Feb 19 02:38:32 2010 KojiConfigurationoplevsIntsant green oplevs for ITMs shooting from the ends

I set up instant green oplevs for ITMs.

A green laser pointer has been set on each end table. It illuminates the ITM center. The beam goea through the ETM substrate.
The reflected green beam returns to the ETM if the ITMs are aligned. Even though the reflected beam to the end is too big, this can
be a rough reference for each ITM.

Note: The green laser pointer at the ETMX were borrowed from Frank. We must return it to him when we finish the work.

  2614   Fri Feb 19 00:31:17 2010 JenneUpdateCOCNew ITMX guiderods glued

[Jenne, Kiwamu, with moral support from Koji, and loads of advice from Steve and Bob]

New upgrade ITMX (ITMU03) has it's guiderod & standoff glued on, as step 1 toward hanging the ITMs.

Procedure:

1. Make sure you have everything ready.  This is long and complicated, but not really worth detail here.  Follow instructions in E970037 (SOS Assembly Spec), and get all the stuff in there.

2. Set optic in a 'ring stand', of which Bob has many, of many different sizes. They are cleaned and baked, and in the cleanroom cupboard on the bottom just behind the door. We used the one for 3" optics.  This lets you sit the optic down, and it only rests on the bevel on the outside, so no coated surface touches anything.

3. Drag wipe the first surface of the optic, using Isopropyl Alcohol.  We used the little syringes that had been cleaned for the Drag Wipe Event which happened in December, and got fresh Iso out of the bottle which was opened in Dec, and put it into a baked glass jar.  The drag wipe procedure was the same as for the December event, except the optic was flat on the bench, in the ring holder.

4. Turn the optic over.

5. Drag wipe the other surface.

6. Align the optic in the guiderod gluing fixture (Step 3 in Section 3.2.1: Applying Guide Rod and Wire Standoff of E970037).

7. Set guiderod and standoff (1 guiderod on one side, 1 standoff on the other, per instructions) against the side of the optic.

8.a.  Use a microscope mounted on a 3-axis micrometer base to help align the guiderod and standoff to the correct places on the optic (Steps 4-5 of Section 3.2.1).  This will be much easier now that we've done it once, but it took a looooooong time. 

8.b.  We put the optic in 180deg from the way we should, based on the direction of the wedge angle in the upgrade table layout (wedge angle stuff used a "Call a Friend" lifeline.  We talked to Koji.) The instructions say to put the guiderod and standoff "above" the scribe lines in the picture on Page 5 of E970037 - the picture has the arms of the fixture crossing over the scribe lines.  However, to make the optic hang correctly, we needed to put the guiderod and standoff below the scribe lines.  This will be true as long as the arrow scribe line (which marks the skinniest part of the optic, and points to the HR side) is closest to you when the optic is in the fixture, the fixture is laying on the table (not standing up on end) with the micrometer parts to your right.  We should put the other ITM into the fixture the other way, so that the arrow is on the far side, and then we'll glue the guiderod and standoff "above" the scribe lines.  Mostly this will be helpful so that we can glue in exactly the places the instructions want us to.

8.c.  The biggest help was getting a flashlight to help illuminate the scribe lines in the optic while trying to site them in the microscope.  If you don't do this, you're pretty much destined to failure, since the lights in the cleanroom aren't all that bright. 

8.d.  The micrometer mount we were able to find for the microscope has a max travel of 0.5", but the optic is ~1" thick.  To find the center of the optic for Step 5 in the guiderod and standoff alignment we had to measure smaller steps, such as bevel-to-end-of-scribe-line, and length-of-scribe-line then end-of-scribe-line-to-other-bevel.  Thankfully once we found the total thickness and calculated the center, we were able to measure once bevel-to-center. 

9. Apply glue to the guiderod and standoff.  We made sure to put this on the "down" side, which once the optic is hung, will be the top of the little rods.  This matches the instructions as to which side of the rods to apply the glue on.  The instructions do want the glue in the center of the rod though, but since we put the optic in the fixture the wrong way, we couldn't reach the center, so we glued the ends of the rods.  We will probably apply another tiny dab of glue on the center of the rod once it's out of the fixture, perhaps while the magnet assemblies are being glued.

10.  We didn't know if the airbake oven which Bob showed us to speed up the curing of our practice epoxy last night was clean enough for the ITM (he was gone by the time we got to that part), so for safety, we're leaving the optic on the flow bench with a foil tent (the foil is secured so there's no way it can blow and touch the optic).  This means that we'll need the full curing time of the epoxy, not half the time.  Maybe tomorrow he'll let us know that the oven is in fact okay, and we can warm it up for the morning.

  2613   Thu Feb 18 15:39:16 2010 Koji and SteveConfigurationVACvalve condition: ALL OFF

As preparation for the upcoming planned power outage we turned turbos, RGA off and closed valves.

IFO chamber is not pumped now. Small leaks and out gassing will push the pressure up slowly. At 3 mTorr of P1 the PSL output shutter

will be closed by the interlock.

It is OK to use light in the IFO up to this point.

  2612   Thu Feb 18 10:10:43 2010 steveConfigurationGeneral480 V AC power turned off

Only the 40m cranes are running on 480VAC The electricians are rewiring this transformer on the mezzanine so it was shut down.

I tested all three cranes before the 480V power was turned off. The last thing to do with the cranes to wipe them down before use.

It will happen on next Tuesday morning.

  2611   Wed Feb 17 19:36:05 2010 KojiUpdateCOCArm visibility

I have measured the arm visibilities.
I did not see any change since the last wiping. Our vacuum is not contaminating the cavity in the time scale of 2 months.

It is very good.


Arm visibility measurement ~ latest (Feb. 17, 2010)

X Arm: 0.898 +/- 0.003
Y Arm: 0.892 +/- 0.006

Arm visibility measurement after the vent (Dec. 14, 2009)

X Arm: 0.897 +/- 0.005
Y Arm: 0.893 +/- 0.004

Arm visibility measurement before the vent (Nov 10, 2009)

X Arm: 0.875 +/- 0.005
Y Arm:
0.869 +/- 0.006

  2610   Wed Feb 17 12:45:19 2010 josephbUpdateComputersUpdated Megatron and its firewall

I updated the IP address on the Cisco Linksys wireless N router, which we're using to keep megatron separated from the rest of the network.  I then went in and updated megatrons resolv.conf and host files.  It is now possible to ssh into megatron again from the control machines.

  2609   Tue Feb 16 16:24:30 2010 kiwamuUpdateGreen LockingRe:Re:take some optics away from the ETM end table

I put the TRY_PD back to the end table and aligned it. Now it seems to be working well.

Quote:

The PD Kiwamu removed from the Y table was TRY, which we still need.

My bad if he took that. By mistake I told him that was the one I installed on the table for the length measurement and we didn't need it anymore.

I'm going to ask Kiwamu if he can kindly put it back.

 I am going to put the PD back to the Y end table in this afternoon.

 

  2608   Tue Feb 16 15:25:00 2010 AlbertoConfigurationLSCArms and PRC not locking

 

 shock.jpg

  2607   Tue Feb 16 14:10:06 2010 josephb, rob, kojiConfigurationLSCArms and PRC not locking

Quote:

Since last Friday either the arms or the PRC can't lock.

The montors show the beam flashing on the end mirrors, but the cavity can't get locked. The error signal looks fine. I suspect a computer problem.

Also PRC can't lock. SPOB is suspiciously stuck at about -95. Although that's not a fixed number, but covering the by hand the SPOB PD on the ITMY table doesn't change the number. I check the DC output of the photodetector and it is actually seen the beam.

Suspecting computer problems started after last Thursday's IP switch, I rebooted the frame builder, c1dcuepics, c1daqctrl and all the front ends. I then burtrestored to February 1st at 1:00 am.

Before I burtrestored c1iscepics, SPOB had gone back to more typical numbers around 0, as it usually read when PRC wasn't locked.

But burtrestoring c1iscepics, return it to the -95 of earlier.

Burterestoring to other times or dates didn't solve the problems.

 Koji and I started poking around, trying to understand what was going on.  At first, we thought it might be related to a computer error, as it seemed.

Fortunately, Rob stopped by and explained that the boost stage of the filter comes under c1lsc control, and will be turned on or off depending on the power in the arms.  Although if you turn it off, it will remain off, it just if its manually selected on, it may go on or off.

Similarly, the output from the Xarm filter bank to the ETMX  filter input will be turned on or off depending on the power in the arm.

Anyways, the locking trouble turns out to be due to no RF sidebands at 33 MHz.  The output of the Marconi was unplugged.  I don't know who, or why did it, but I've plugged it in for now, so we can lock the arms.  Let us know if you need in unplugged.  Thanks.

  2606   Tue Feb 16 11:12:51 2010 kiwamuUpdateGreen LockingRe:take some optics away from the ETM end table

Quote:

Quote:

In the last two days Steve and I took some optics away from the both ETM end table.

This is because we need an enough space to set up the green locking stuff into the end table, and also need to know how much space is available.

Optics we took away are : Alberto's RF stuff, fiber stuff and some optics obviously not in used.

The picture taken after the removing is attached. Attachment1:ETMX, Attachment2:ETMY

And the pictures taken before the removing are on the wiki, so you can check how they are changed.

http://lhocds.ligo-wa.caltech.edu:8000/40m/Optical_Tables

The PD Kiwamu removed from the Y table was TRY, which we still need.

My bad if he took that. By mistake I told him that was the one I installed on the table for the length measurement and we didn't need it anymore.

I'm going to ask Kiwamu if he can kindly put it back.

 I am going to put the PD back to the Y end table in this afternoon.

  2605   Tue Feb 16 10:01:16 2010 AlbertoConfigurationLSCArms and PRC not locking

Since last Friday either the arms or the PRC can't lock.

The montors show the beam flashing on the end mirrors, but the cavity can't get locked. The error signal looks fine. I suspect a computer problem.

Also PRC can't lock. SPOB is suspiciously stuck at about -95. Although that's not a fixed number, but covering the by hand the SPOB PD on the ITMY table doesn't change the number. I check the DC output of the photodetector and it is actually seen the beam.

Suspecting computer problems started after last Thursday's IP switch, I rebooted the frame builder, c1dcuepics, c1daqctrl and all the front ends. I then burtrestored to February 1st at 1:00 am.

Before I burtrestored c1iscepics, SPOB had gone back to more typical numbers around 0, as it usually read when PRC wasn't locked.

But burtrestoring c1iscepics, return it to the -95 of earlier.

Burterestoring to other times or dates didn't solve the problems.

  2604   Tue Feb 16 09:51:22 2010 AlbertoUpdateGreen Lockingtake some optics away from the ETM end table

Quote:

In the last two days Steve and I took some optics away from the both ETM end table.

This is because we need an enough space to set up the green locking stuff into the end table, and also need to know how much space is available.

Optics we took away are : Alberto's RF stuff, fiber stuff and some optics obviously not in used.

The picture taken after the removing is attached. Attachment1:ETMX, Attachment2:ETMY

And the pictures taken before the removing are on the wiki, so you can check how they are changed.

http://lhocds.ligo-wa.caltech.edu:8000/40m/Optical_Tables

The PD Kiwamu removed from the Y table was TRY, which we still need.

My bad if he took that. By mistake I told him that was the one I installed on the table for the length measurement and we didn't need it anymore.

I'm going to ask Kiwamu if he can kindly put it back.

  2603   Sat Feb 13 18:58:31 2010 josephb, alexUpdateComputersfb40m testpoints fixed

I received an e-mail from Alex indicating he found the testpoint problem and fixed it today:

Quote from Alex: "After we swapped the frame builder computer it has reconfigured all device files and I needed to create some symlinks on /dev/ to make tpman work again. I test the testpoints and they do work now."

 

  2602   Sat Feb 13 13:21:53 2010 KojiUpdateElectronicstriple resonant EOM --- liniaryity test

Looks good. I just thought of the idea that we also can use Alberto's PLL setup to sense the modulation with more sensitivity.  ;-)

Quote:

I have measured the linearity of our triple resonant EOM (i.e. modulation depth versus applied voltage)

The attached figure is the measured modulation depth as a function of the applied voltage.

The linear behavior is shown below 4Vrms, this is good.

Then it is  slowly saturated as the applied voltage goes up above 4Vrms.

However for the resonance of 29.5MHz, it is difficult to measure below 7Vrms because of the small modulation depth.

Our triple resonant EOM looks healthy

 - - - - result from fitting - - -

11MHz: 910mrad/Vrms+20mrad

29.5MHz: 46mrad/Vrms+6.2mrad

55MHz:820mrad/Vrms+10mrad

 

  2601   Fri Feb 12 18:58:46 2010 kiwamuUpdateGreen Lockingtake some optics away from the ETM end table

In the last two days Steve and I took some optics away from the both ETM end table.

This is because we need an enough space to set up the green locking stuff into the end table, and also need to know how much space is available.

Optics we took away are : Alberto's RF stuff, fiber stuff and some optics obviously not in used.

The picture taken after the removing is attached. Attachment1:ETMX, Attachment2:ETMY

And the pictures taken before the removing are on the wiki, so you can check how they are changed.

http://lhocds.ligo-wa.caltech.edu:8000/40m/Optical_Tables

Attachment 1: DSC_1164.JPG
DSC_1164.JPG
Attachment 2: DSC_1172.JPG
DSC_1172.JPG
  2599   Fri Feb 12 15:59:16 2010 josephbUpdateComputersTestpoints not working

Non-testpoint channels seem to be working in data viewer, however testpoints are not.  The tpman process is not running on fb40m.  My rudimentary attempts to start it have failed. 

# /usr/controls/tpman &
13929
# VMIC RFM 5565 (0) found, mapped at 0x2868c90
VMIC RFM 5579 (1) found, mapped at 0x2868c90
Could not open 5565 reflective memory in /dev/daqd-rfm1
16 kHz system
Spawn testpoint manager
no test point service registered
Test point manager startup failed; -1

It looks like it may be an issue with the reflected memory (although the cables are plugged in and I see the correct lights lit on the RFM card in back of fb40m.)

The fact that this is a RFM error is confirmed by /usr/install/rfm2g_solaris/vmipci/sw-rfm2g-abc-005/util/diag/rfm2g_util and entering 3 (which should be the device number).

Interestingly, the device number 4 works, and appears to be the correct RFM network (i.e. changing ETMY lscPos offset changes to the corresponding value in memory).

So, my theory is that when Alex put the cards back in, the device number (PCI slot location?) was changed, and now the tpman code doesn't know where to look for it.

Edit: Doesn't look like PCI slot location is it, given there's 4 slots and its in #3 currently (or 2 I suppose, depending on which way you count).  Neither seems much the number 4.  So I don't know how that device number gets set.

 

 

  2598   Fri Feb 12 14:19:28 2010 rana, steveHowToloreInternational Fax

Steve showed me how to send an international fax today:

  1. Load paper.
  2. Dial:   011 - (country code) - number
  3. Press START (either the black or color option)
  4. wait for the screaming fax noise
  5. Done

 

  2597   Fri Feb 12 13:56:16 2010 josephbUpdateComputersFinishing touches on IP switch over

The GPIB interfaces have been updated to the new 192.168.113.xxx addresses, with Alberto's help.

Spare ethernet cables have been moved into a cabinet halfway down the x-arm.

The illuminators have a white V error on the alarm handler, but I'm not sure why.  I can turn them on and off using the video screen controls (except for the x arm, which has no computer control, just walk out and turn it on).

There's a laptop or two I haven't tracked down yet, but that should be it on IPs. 

At some point, a find and replace on 131.215.xxx.yyy addresses to 192.168.xxx.yyy should be done on the wiki.  I also need to generate an up to date ethernet IP spreadsheet and post it to the wiki.

 

  2596   Fri Feb 12 13:15:41 2010 kiwamuUpdateElectronicstriple resonant EOM --- liniaryity test

I have measured the linearity of our triple resonant EOM (i.e. modulation depth versus applied voltage)

The attached figure is the measured modulation depth as a function of the applied voltage.

The linear behavior is shown below 4Vrms, this is good.

Then it is  slowly saturated as the applied voltage goes up above 4Vrms.

However for the resonance of 29.5MHz, it is difficult to measure below 7Vrms because of the small modulation depth.

Our triple resonant EOM looks healthy

 - - - - result from fitting - - -

11MHz: 91mrad/Vrms+2.0mrad

29.5MHz: 4.6mrad/Vrms+6.2mrad

55MHz:82mrad/Vrms+1.0mrad

Attachment 1: linearity_edit.png
linearity_edit.png
  2595   Fri Feb 12 11:56:02 2010 josephbUpdateComputersNodus slow ssh fixed

Koji pointed out that logging into Nodus was being abnormally slow.  I tracked it down to the fact we had forgotten to update the address for the DNS server running on linux1 in the resolv.conf file on nodus.  Basically it was looking for a DNS server which didn't exit, and thus was timing out before going to the next one.  SSHing into nodus should be more responsive.

  2594   Fri Feb 12 11:44:11 2010 josephbUpdateComputersStatus of the IP change over

Quote:

After Joe left:

  1. Turned on op440m and returned him his keyboard and mouse.
  2. Damped MC2.
  3. Opened PSL shutter - locked PMC, FSS,
  4. Started StripTool displays on op540m.
  5. op340m doesn't respond to ping from anyone.
  6. started FSS  SLOW and RCPID scripts on op540 - need to kill and restart on op430m.
  7. ASS wouldn't come up - it doesn't know who linux1 is.
  8. MC autolocker wouldn't run on op540m because of a perl module issue, started it on op440m - it needs to be killed and restarted on op430m.
  9. probably mafalda, linux2, and op430m need some attention - they are all in the same rack.

As of 7:18 PM, the MC is locked and the PSL seems normal + all suspensions are damped and the ELOG is back up as well as the SVN.

5) op340m has had its hosts table and other network files updated.  I also removed its outdated hosts.deny file which was causing some issues with ssh.

6) On op340m I started FSSSlowServo, with "nohup ./FSSSlowServo", after killing it on op540m.

I also kill RCthermalPID.pl, and started with "nohup ./RCthermalPID.pl" on op540m.

7) c1ass is fixed now.  There was a typo in the resolv.conf file (namerserver -> nameserver) which has been fixed.  It is now using the DNS server running on linux1 for all its host name needs.

8) I killed the autlockMCmain40m process running on op440m, modified the script to run on op340m, logged into op340m, went to /cvs/cds/caltech/scripts/MC and ran nohup ./autolockMCmain40m

9) Linux2 does not look like it has not been connected for awhile and its wasn't connected when we started the IP change over yesterday.  Is it supposed to still be in use?  If so, I can hook it up fairly easily.  op340m, as noted earlier, has been switched over.  Mafalda has been switched over.

10) c0rga has now been switched over. 

11) aldabella, the vacuum laptop has had its starting environment variables tweaked (in the /home/controls/.cshrc file) so that it looks on the 192.168.113 network instead of the 131.215.113.  This should mean Steve will not have any more trouble starting up his vacuum control screen.

12) Ottavia has been switched over.

13) At this time, only the GPIB devices and a few laptops remain to get switched over.

  2593   Thu Feb 11 19:20:44 2010 ranaUpdateComputersStatus of the IP change over

After Joe left:

  1. Turned on op440m and returned him his keyboard and mouse.
  2. Damped MC2.
  3. Opened PSL shutter - locked PMC, FSS,
  4. Started StripTool displays on op540m.
  5. op340m doesn't respond to ping from anyone.
  6. started FSS  SLOW and RCPID scripts on op540 - need to kill and restart on op430m.
  7. ASS wouldn't come up - it doesn't know who linux1 is.
  8. MC autolocker wouldn't run on op540m because of a perl module issue, started it on op440m - it needs to be killed and restarted on op430m.
  9. probably mafalda, linux2, and op430m need some attention - they are all in the same rack.

As of 7:18 PM, the MC is locked and the PSL seems normal + all suspensions are damped and the ELOG is back up as well as the SVN.

  2592   Thu Feb 11 18:53:57 2010 steveConfigurationVACVAC NORMAL is back

Quote:

Joe and Alex are working on the computers. Our vacuum system is temporary "All off" condition: meaning all valves are closed, so there is no pumping. cc1 = 1.6e-6 Torr

 Designated vacuum control lap top is trouble some to use. Joe finally fixed it and I switched valve configuration back to vacuum normal. Shutter is open

  2591   Thu Feb 11 18:33:54 2010 josephb, alexUpdateComputersStatus of the IP change over

A few machines have still not been changed over, including a few laptops, mafalda, ottavia, and c0rga.

All the front ends have been changed over.

fb40m died during a reboot and was replaced with a spare Sun blade 1000 that Larry had.  We had to swap in our old hard drive and memory.

All the front ends, belladonna, aldabella, and the control room machines have been switched over. Nodus was changed over after we realized we hosed the elog and svn by switching linux1's IP.

At this point, 90% of the machines seem to be working, although c0daqawg seems to be having some issues with its startup.cmd code.

  2590   Thu Feb 11 16:52:53 2010 kiwamuUpdateElectronicstriple resonant EOM ---- preliminary result

The commercial resonant EOM of New Focus has the modulation efficiency of 50-150mrad/Vrms. ( This number is only true for those EOM made from KTP such as model4063 and model4463 )

Our triple-resonant EOM (made from KTP as well) has a 90mrad/Vrms and 80mrad/Vrms at the reosonances of 11MHz and 55MHz respectively.

Therefore our EOM is as good as those of company-made so that we can establish a new EOM company

Quote:

Hey, this looks nice, but can you provide us the comparison of rad/V with the resonant EOM of New Focus?

Quote:

I have made a prototype circuit of the triple resonant EOM.

The attached is the measured optical response of the system.

The measured gains at the resonances are 8.6, 0.6 and 7.7 for 11MHz, 29.5MHz and 55MHz respectively.

I successfully got nice peaks at 11MHz and 55MHz. In addition resultant optical response is well matched with the predicted curve from the measured impedance.

However there is a difference from calculated response (see past entry) (i.e. more gains were expected)

Especially for the resonance of 29.5MHz, it was calculated to have gain of 10, however it's now 0.6. Therefore there must a big loss electrically around 29.5MHz.

I am going to re-analyze the impedance and model the performance in order to see what is going on.

 

 

  2589   Thu Feb 11 15:53:59 2010 steveConfigurationVACPSL output shutter is closed

Joe and Alex are working on the computers. Our vacuum system is temporary "All off" condition: meaning all valves are closed, so there is no pumping. cc1 = 1.6e-6 Torr

  2588   Wed Feb 10 23:44:56 2010 KojiSummaryCOCPhase Map Analysis

In the middle of the last month, Kiwamu and I went to Garilynn's lab to measure the phase maps of the new ITMs and SRMs.

Analysis of the phase map data were posted on the svn directory:
https://nodus.ligo.caltech.edu:30889/svn/trunk/docs/upgrade08/cocdocs/PhaseMaps/

The screen shots and the plots were summarized in a PDF file. You can find it here:
http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/Main_Optics_Phase_Maps

The RoCs for all of the PRMs are turned out to be ~155m. This is out of the spec (142m+/-5m) although the actual effect is not understand well yet..

These RoCs are almost independent from the radus of the assumed gaussian beam.
In deed, I have checked the dependence of the RoC on the beam spot position, and it turned out that the RoCs vary only little.
(In the SRMU01 case, for example, it varies from 153.5m to 154.9m.)
The beam radius of 3mm was assumed. The RoCs were calculated 20x20mm region around the center of the mirror with a 2mm mesh.
 

Attachment 1: SRM01_HR_RoC_rad_15mm.png
SRM01_HR_RoC_rad_15mm.png
Attachment 2: SRM01_HR_RoC_scan.png
SRM01_HR_RoC_scan.png
ELOG V3.1.3-